# AcousTools: A ‘Full-Stack’, Python-Based, Acoustic Holography Library

### v3

@@ Abstract
preprint: AIP/123-QED
@@
I Introduction
Non-contact manipulation and stimulus open the door for numerous applications in a user-friendly and ergonomic way. Objects can be fabricated without ever touching the materials used, either by assembling smaller building blocks LeviPrint or by shaping raw materials AcoustoFab. Haptic stimulus can give greater immersion in virtual reality HapticVR; MidAirKeyboard, leveraging human tactile mechanisms to produce higher quality stimulus HapticsMPSTM; HapticIllusions and render complex stimulus such as softness HapticSoftness. Levitation of small particles, in addition to fabrication, allows the creation of computer displays that are truly 3D VolDisplaySingleParticle, by using props to represent objects Articulev or by exploiting persistence of vision (PoV) to render shapes in 3D GSPAT; VolDisplaySingleParticle. All these applications and more are unlocked by utilising focused mid-air ultrasound.
To create these applications, a software package or library that can model the behaviour of acoustic waves is advantageous OpenMPD; Ultraino. Common tasks include computing phases for transducer arrays in complex scenarios LeviPrint, predicting how sound scatters from irregular objects HighSpeedBEM, and analysing the force exerted on an object in a sound field AcousticElements. Pre-built software allows developers to focus on higher-level goals without needing to implement these complex processes from scratch.
@@ I Introduction
There are a variety of software packages that attempt to meet the stages described in the full-stack OpenMPD; Ultraino; Levitate; Kwave; FOCUS. However, all existing solutions fail to meet one or many of these categories as can be seen in Tab.1. As a result, developers often need to switch between different packages or are unable to use certain methodologies without implementing them independently.
@@
In this work, we step through each stage of the full-stack, showing AcousTools’ efficacy in that area and highlighting where other software falls short. Its implementation using PyTorch in Python leverages the language’s ease of access especially for beginners pythonEase, as well as the parallelised and comprehensive capabilities of the PyTorch environment PyTorch. AcousTools is therefore the general purpose software solution that the acoustic holography community currently lacks. We anticipate that the community will adopt it as a standard code library, leading to a greater degree of interoperability and reproducibility among researchers.
@@
II Related Work
II.1 Acoustic Holography
@@
Acoustic holography, used to produce 3D sound fieldsHoloForAcoustics; AcousticElements; GSPAT; HighSpeedBEM, can be used to focus high frequency sound waves. This allows for non-contactless manipulation VolDisplaySingleParticle; LeviPrint and stimulation eigensolver; HapticsMPSTM of materials at a distance. Unlike magnetic levitation, it is not limited to ferrous materials, and unlike optical holography using lasers, there is no dangerous heating of the samples 2DDroplets. Thus, acoustic holography is an exciting and interactive method. Acoustic levitation is even used in the field of nanoparticle mid-air chemistry AcousticChemistry, but only the simplest acoustic setups TinyLev are used in these cases due to the complexity of the control software commonly used in holographic work. Modern open-source hardwareUltraino; OpenMPD combined with reliable and flexible software has a potential to expand the domains where acoustic holography can be used.
@@
Propagation Methods: Acoustic holography computes a set of transducer activations used to produce a desired 3D sound field AcousticElements. It predicts how these waves interfere and build up VolDisplaySingleParticle; GSPAT; Diff-PAT by modelling sound propagation. Sound waves are emitted from a source, often an array of transducers, and typically modelled using the Piston Model GSPAT; LeviPrint; HapticsMPSTM; VolDisplaySingleParticle; AcoustoFab. In this model, the propagation from a single transducer $t\in\bm{\tau}$, for a set of transducers $\bm{\tau}$, located at $\bm{z}_{t}$ to a point in space located at $\bm{z}_{n}$ can be modelled as GSPAT
$F_{n,t}=\frac{2p_{ref}J_{1}(kr\sin(\theta))}{kr\sin(\theta)d(\bm{z}_{n},\bm{z}_{t})}e^{ikd(\bm{z}_{n},\bm{z}_{t})}.$ (1)
@@
Here, $p_{ref}$ is the reference pressure at 1m for the transducer, $\theta$ is the angle between the transducer normal and the point, $r$ is the transducer radius, $k$ is the wavenumber, $J_{1}$ is the Bessel function of the first kind, and $d$ is the distance from the transducer and the point. The actual complex pressure can be then found by multiplying $F_{n,t}$ by the activation of that transducer, $x_{t}=A_{t}e^{i\varphi_{t}}$ where $A_{t}$ and $\varphi_{t}$ are the amplitude and phase, respectively. The values of $F_{n,t}$ and $x_{t}$ for each transducer and point can be combined into matrix forms, and the overall acoustic pressure field at a set of $N$ points, $\bm{z}_{n}\in\bm{\nu},|\bm{\nu}|=N$, is given by
$\bm{p}_{\bm{\nu}}=\bm{F}_{\bm{\nu}\bm{\tau}}\ \bm{x}_{\bm{\tau}}.$ (2)
@@
In a free-field scenario without scattering or occlusion, this model is sufficient. However, considering scattering would provide benefits to many applications. When there are objects in a levitation domain, for example when an object is fabricated during levitation, modelling sound scattering can increase manipulation precision and reduce potential levitation failure AcoustoFab. This can be done by using a three step model, known as the boundary element method (BEM). Instead of the free-field propagator $\bm{F}_{\bm{\nu}\bm{\tau}}$, this model uses an extended propagator $\bm{E}_{\bm{\nu}\bm{\tau}}$, which is formed as HighSpeedBEM
$\bm{E}_{\bm{\nu}\bm{\tau}}=\bm{F}_{\bm{\nu}\bm{\tau}}+\bm{G}_{\bm{\nu}\bm{\mu}}\bm{H}_{\bm{\mu}\bm{\tau}},$ (3)
@@
Where $\bm{F}_{\bm{\nu}\bm{\tau}}$ is the same as in Eq.1, $\bm{H}_{\bm{\mu}\bm{\tau}}$ models the propagation of the sound from the transducers to a mesh, $\bm{\mu}$, representing a scattering object, and $\bm{G}_{\bm{\nu}\bm{\mu}}$ models how the waves travel from the mesh surface to the points. Thus, the overall propagator $\bm{E}_{\bm{\nu}\bm{\tau}}$ considers both the direct ($\bm{F}_{\bm{\nu}\bm{\tau}}$) and the scattered ($\bm{G}_{\bm{\nu}\bm{\mu}}\bm{H}_{\bm{\mu}\bm{\tau}}$) components. However, while using BEM improves many applicationsHighSpeedBEM; ToBEMorNot, the increased computational complexity, especially in computing $\bm{H}_{\bm{\mu}\bm{\tau}}$ holds back its utility ToBEMorNot. As $\bm{H}_{\bm{\mu}\bm{\tau}}$ depends only on the geometry of the scattering objects, it can be precomputed and cached to reduce this cost in some cases HighSpeedBEM. However, if the object’s geometry changes, $\bm{H}_{\bm{\mu}\bm{\tau}}$ must be recomputed, and so dynamic scenarios will suffer from the increase in computation speedsToBEMorNot.
@@
In this paper, we will use $\bm{A}$ to denote an arbitrary propagator, instead of the specific $\bm{F}_{\bm{\nu}\bm{\tau}}$, $\bm{E}_{\bm{\nu}\bm{\tau}}$ or any other propagator. The index, $\bm{\nu}\bm{\tau}$, will be omitted but any propagator will be assumed to have this same form, propagating from a transducer array to a set of points. In the same way, $\bm{x}$ will refer to a generic set of transducer activations with the $\bm{\tau}$ index omitted.
@@
Phase Retrieval Algorithms: For a given propagator, the transducer activations $\bm{x}$, which we refer to as an acoustic hologram, need to be determined. A generalised objective function can be defined as TGS
@@
Here, $\mathcal{L}$ is an objective function of the complex field produced by the propagation of the hologram x, for example maximising the pressure amplitude at the targetsIB; GSPAT; TGS. The constraint $\mathcal{C}$ quantifies hardware requirements for the hologram x, often using indicator functions that are zero if these requirements are met and infinite otherwise. Common constraints include forcing transducer amplitudes to be maximum IB, $|\textbf{{x}}|=1$, or allowing amplitudes to vary within physically feasible limits GSPAT, $|\textbf{{x}}|\leq 1$. Methods to fulfil Eq.4 take two main forms, gradient-descent as well as projective methods.
@@
In order to then perform this optimisation, computing the gradients can be done either explicitly or implicitly. Explicit computations require more labour from the developer AcousticElements, where they must directly know and implement how to find the gradients. Conversely, implicit gradients are more convenient and versatile Diff-PAT; Diff-PAT-Expanded as the exact form of the gradients are avoided by utilising automatic differentiation. This accumulates the gradients through computation of an objective and then uses the chain rule to backpropagate through this computation tree to compute the gradients of a wide variety of functions Diff-PAT; Diff-PAT-Expanded. This gives flexibility in algorithm design as the gradient to minimise an objective need not to be defined directly. However, to use automatic differentiation, the chain of operations needs to be maintained, requiring specific implementations such as that given in the Python packages PyTorchPyTorch or Jax jax.
On the other hand, alternating projections are based on – normally iterative – algorithms, in which forward and backward propagations between the ultrasound source and the target points are combined with projections to feasible values. The most common algorithms are based on the Gerchberg-Saxton (GS) algorithmGS, first described in acoustics as the iterative backpropagation algorithm (IB), which can be written as
$\textit{{x}}_{k+1}^{IB}=\mathcal{P}_{\chi}(\textit{{A}}^{*}(\mathcal{P}_{\Pi}(\textit{{A}}\textit{{x}}_{k}|\bm{y}))).$ (6)
In this case, $\mathcal{P}_{\Pi}$ is introduced as a second indicator constraint that sets the point amplitudes to some target point amplitude y, defined as $\mathcal{P}_{\Pi}(\textbf{{p}}|\bm{y})=(\textbf{{p}}/|\textbf{{p}}|)\cdot\bm{y}$ given the point activation p and element wise multiplication $(\cdot)$. $\bm{A}$ propagates the hologram to the target points pressure, this is then constrained by $\mathcal{P}_{\Pi}$. This constrained point is back-propagated using the conjugate transpose of the propagator, $\bm{A}^{*}$, which is itself constrained by $\mathcal{P}_{\chi}$. The IB algorithm is known to act as projected gradient descent for an objective function that maximizes the sum of amplitudes at target pointsTGS.
@@
$\bm{x}^{naive}=\mathcal{P}_{\chi}(\bm{A}^{*}\bm{y}).$ (7)
This results in very fast computation rates, but failing to finely equalize pressure at target points. Alternatively, GS-PAT removes the target points amplitude constraints, combining forward and backward propagation for high computational speed GSPAT, while Weighted Gerchberg-Saxton (WGS) weights the pressure at each target point during each iteration to achieve low variance in pressure across target pointsTGS. Nevertheless, beyond interpretability, alternating projection algorithms also lack gradient descent’s adaptability to adapt to custom objectives that may be required by users for different applications.
Gor’kov potential, Force, and Stiffness: When the field has been modelled for a given set of transducer activations, various metrics can be computed that quantify the field. The simplest measure of an acoustic field is the pressure at a point in space. When the solver has correctly focused waves at a point, there should be a local maxima of pressure at that point and the amplitude of the sound field at position can be described by $|\bm{A}\bm{x}|$. Similarly, the phase can be found by taking the complex argument of $\bm{A}\bm{x}$.
The Gor’kov potential is a potential field based on both the pressure and gradient of pressure surrounding a small particle as AcousticElements
@@
Here, $c_{0}$ and $c_{p}$ are the speed of sound in the medium and the particle, $\rho_{0}$ and $\rho_{p}$ are the density of the medium and particle, $\omega$ is the angular frequency, and $V$ is the volume of the particle. $p$ is the pressure at a point and ${\partial p}/{\partial a}$ is the gradient of pressure in the $a$-axis where $a\in(x,y,z)$ AcousticElements. The Gor’kov potential is the field such that the negative gradient of the Gor’kov potential is equal to the force acting on a small particle in the sound field.
$\bm{F}_{rad}=-\bm{\nabla}U.$ (11)
As a result, the point at which a particle will levitate corresponds to where the gradient of Gor’kov potential is zero (so there is no net force at that point). In reality, the levitation location will be very slightly below this point due to the force of gravity but this difference can be largely neglected for very small particles ReviewOfProgress. The points with $\bm{\nabla}U=0$ includes both minima and maxima of Gor’kov potential and both are technically valid levitation locations (with net force equal to zero). However, at a maxima of Gor’kov potential, the forces diverge, and therefore the trap becomes unstable. To quantify a stable trap, we can use the stiffness at a point in space, which can be represented as the sum of the unmixed, second gradients (the Laplacian) of the Gor’kov potential as AcousticElements
@@
II.2 Existing Software
AcousTools OpenMPDOpenMPD UltrainoUltraino LevitateLevitate

### v6

@@ Abstract
{IEEEkeywords} Acoustic applications, Acoustic field, Levitation, Software libraries
@@
1 Introduction
Highlights
$\bullet$ Acoustic holography has neither a unified framework to think about acoustic applications nor a single software library to implement them – reducing interoperability and flexibility.
$\bullet$ We present the Acoustic Full-Stack, which describes the process a developer goes through to implement an acoustic application and then present AcousTools; a python-based library to implement each of these stages.
$\bullet$ With a unified software library, the interpretability between researchers will increase and the full-suite of options means developers can make fully informed choices about which methods and algorithms to use.
[…]
Non-contact manipulation and stimulus open the door for numerous applications in a user-friendly and ergonomic way. Objects can be fabricated without ever touching the materials used, either by assembling smaller building blocks [LeviPrint] or by shaping raw materials [AcoustoFab]. Haptic stimulus can give greater immersion in virtual reality [HapticVR], leveraging human tactile mechanisms to produce higher quality stimulus [HapticsMPSTM, HapticIllusions] and render complex stimulus such as softness [HapticSoftness]. Levitation of small particles, in addition to fabrication, allows the creation of computer displays that are truly 3D [VolDisplaySingleParticle], by using props to represent objects [Articulev] or by exploiting persistence of vision (PoV) to render shapes in 3D [GSPAT, VolDisplaySingleParticle]. All these applications and more are unlocked by utilising focused mid-air ultrasound.
To create these applications, a software package or library that can model the behaviour of acoustic waves is advantageous [OpenMPD, Ultraino]. Common tasks include computing phases for transducer arrays in complex scenarios [LeviPrint], predicting how sound scatters from irregular objects [HighSpeedBEM], and analysing the force exerted on an object in a sound field [AcousticElements]. Pre-built software allows developers to focus on higher-level goals without needing to implement these complex processes from scratch.
@@ 1 Introduction
There are a variety of software packages that attempt to meet the stages described in the full-stack [OpenMPD, Ultraino, Levitate, Kwave, FOCUS]. However, all existing solutions fail to meet one or many of these categories as can be seen in Tab.1. As a result, developers often need to switch between different packages or are unable to use certain methodologies without implementing them independently.
@@
In this work, we step through each stage of the full-stack, showing AcousTools’ efficacy in that area and highlighting where other software falls short. Its implementation using PyTorch in Python leverages the language’s ease of access especially for beginners [pythonEase], as well as the parallelised and comprehensive capabilities of the PyTorch environment [PyTorch]. AcousTools is therefore the general purpose software solution that the acoustic holography community currently lacks. We anticipate that the community will adopt it as a standard code library, leading to a greater degree of interoperability and reproducibility among researchers.
@@
2 Related Work
2.1 Acoustic Holography
@@
Acoustic holography, used to produce 3D sound fields[HoloForAcoustics, AcousticElements, GSPAT, HighSpeedBEM], can be used to focus high frequency sound waves. This allows for non-contactless manipulation [VolDisplaySingleParticle, LeviPrint] and stimulation [eigensolver, HapticsMPSTM] of materials at a distance. Unlike magnetic levitation, it is not limited to ferrous materials, and unlike optical holography using lasers, there is no dangerous heating of the samples [2DDroplets]. Thus, acoustic holography is an exciting and interactive method. Acoustic levitation is even used in the field of nanoparticle mid-air chemistry [AcousticChemistry], but only the simplest acoustic setups [TinyLev] are used in these cases due to the complexity of the control software commonly used in holographic work. Modern open-source hardware[Ultraino, OpenMPD] combined with reliable and flexible software has a potential to expand the domains where acoustic holography can be used.
@@
Propagation Methods: Acoustic holography computes a set of transducer activations used to produce a desired 3D sound field [AcousticElements]. It predicts how these waves interfere and build up [VolDisplaySingleParticle, GSPAT, Diff-PAT] by modelling sound propagation. Sound waves are emitted from a source, often an array of transducers, and typically modelled using the Piston Model [GSPAT, LeviPrint, HapticsMPSTM, VolDisplaySingleParticle, AcoustoFab]. In this model, the propagation from a single transducer $t\in\boldsymbol{\tau}$, for a set of transducers $\boldsymbol{\tau}$, located at $\boldsymbol{z}_{t}$ to a point in space located at $\boldsymbol{z}_{n}$ can be modelled as [GSPAT]
$F_{n,t}=\frac{2p_{ref}J_{1}(kr\sin(\theta))}{kr\sin(\theta)d(\boldsymbol{z}_{n},\boldsymbol{z}_{t})}e^{ikd(\boldsymbol{z}_{n},\boldsymbol{z}_{t})}.$ (1)
@@
Here, $p_{ref}$ is the reference pressure at 1m for the transducer, $\theta$ is the angle between the transducer normal and the point, $r$ is the transducer radius, $k$ is the wavenumber, $J_{1}$ is the Bessel function of the first kind, and $d$ is the distance from the transducer and the point. The actual complex pressure can be then found by multiplying $F_{n,t}$ by the activation of that transducer, $x_{t}=A_{t}e^{i\varphi_{t}}$ where $A_{t}$ and $\varphi_{t}$ are the amplitude and phase, respectively. The values of $F_{n,t}$ and $x_{t}$ for each transducer and point can be combined into matrix forms, and the overall acoustic pressure field at a set of $N$ points, $\boldsymbol{z}_{n}\in\boldsymbol{\nu},|\boldsymbol{\nu}|=N$, is given by
$\boldsymbol{p}_{\boldsymbol{\nu}}=\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}\ \boldsymbol{x}_{\boldsymbol{\tau}}.$ (2)
@@
In a free-field scenario without scattering or occlusion, this model is sufficient. However, considering scattering would provide benefits to many applications. When there are objects in a levitation domain, for example when an object is fabricated during levitation, modelling sound scattering can increase manipulation precision and reduce potential levitation failure [AcoustoFab]. This can be done by using a three step model, known as the boundary element method (BEM). Instead of the free-field propagator $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, this model uses an extended propagator $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, which is formed as [HighSpeedBEM]
$\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}=\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}+\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}},$ (3)
@@
Where $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ is the same as in (1), $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ models the propagation of the sound from the transducers to a mesh, $\boldsymbol{\mu}$, representing a scattering object, and $\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}$ models how the waves travel from the mesh surface to the points. Thus, the overall propagator $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ considers both the direct ($\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$) and the scattered ($\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$) components. However, while using BEM improves many applications[HighSpeedBEM, ToBEMorNot], the increased computational complexity, especially in computing $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ holds back its utility [ToBEMorNot]. As $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ depends only on the geometry of the scattering objects, it can be precomputed and cached to reduce this cost in some cases [HighSpeedBEM]. However, if the object’s geometry changes, $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ must be recomputed, and so dynamic scenarios will suffer from the increase in computation speeds[ToBEMorNot].
@@
In this paper, we will use $\boldsymbol{A}$ to denote an arbitrary propagator, instead of the specific $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ or any other propagator. The index, $\boldsymbol{\nu}\boldsymbol{\tau}$, will be omitted but any propagator will be assumed to have this same form, propagating from a transducer array to a set of points. In the same way, $\boldsymbol{x}$ will refer to a generic set of transducer activations with the $\boldsymbol{\tau}$ index omitted.
@@
Phase Retrieval Algorithms: For a given propagator, the transducer activations $\boldsymbol{x}$, which we refer to as an acoustic hologram, need to be determined. A generalised objective function can be defined as [TGS]
@@
Here, $\mathcal{L}$ is an objective function of the complex field produced by the propagation of the hologram x, for example maximising the pressure amplitude at the targets[IB, GSPAT, TGS]. The constraint $\mathcal{C}$ quantifies hardware requirements for the hologram x, often using indicator functions that are zero if these requirements are met and infinite otherwise. Common constraints include forcing transducer amplitudes to be maximum [IB], $|\textbf{{x}}|=1$, or allowing amplitudes to vary within physically feasible limits [GSPAT], $|\textbf{{x}}|\leq 1$. Methods to fulfil (4) take two main forms, gradient-descent as well as projective methods.
@@
In order to then perform this optimisation, computing the gradients can be done either explicitly or implicitly. Explicit computations require more labour from the developer [AcousticElements], where they must directly know and implement how to find the gradients. Conversely, implicit gradients are more convenient and versatile [Diff-PAT, Diff-PAT-Expanded] as the exact form of the gradients are avoided by utilising automatic differentiation. This accumulates the gradients through computation of an objective and then uses the chain rule to backpropagate through this computation tree to compute the gradients of a wide variety of functions [Diff-PAT, Diff-PAT-Expanded]. This gives flexibility in algorithm design as the gradient to minimise an objective need not to be defined directly. However, to use automatic differentiation, the chain of operations needs to be maintained, requiring specific implementations such as that given in the Python packages PyTorch[PyTorch] or Jax [jax].
On the other hand, alternating projections are based on – normally iterative – algorithms, in which forward and backward propagations between the ultrasound source and the target points are combined with projections to feasible values. The most common algorithms are based on the Gerchberg-Saxton (GS) algorithm[GS], first described in acoustics as the iterative backpropagation algorithm (IB), which can be written as
$\textit{{x}}_{k+1}^{IB}=\mathcal{P}_{\chi}(\textit{{A}}^{*}(\mathcal{P}_{\Pi}(\textit{{A}}\textit{{x}}_{k}|\boldsymbol{y}))).$ (6)
In this case, $\mathcal{P}_{\Pi}$ is introduced as a second indicator constraint that sets the point amplitudes to some target point amplitude y, defined as $\mathcal{P}_{\Pi}(\textbf{{p}}|\boldsymbol{y})=(\textbf{{p}}/|\textbf{{p}}|)\cdot\boldsymbol{y}$ given the point activation p and element wise multiplication $(\cdot)$. $\boldsymbol{A}$ propagates the hologram to the target points pressure, this is then constrained by $\mathcal{P}_{\Pi}$. This constrained point is back-propagated using the conjugate transpose of the propagator, $\boldsymbol{A}^{*}$, which is itself constrained by $\mathcal{P}_{\chi}$. The IB algorithm is known to act as projected gradient descent for an objective function that maximizes the sum of amplitudes at target points[TGS].
@@
$\boldsymbol{x}^{naive}=\mathcal{P}_{\chi}(\boldsymbol{A}^{*}\boldsymbol{y}).$ (7)
This results in very fast computation rates, but failing to finely equalize pressure at target points. Alternatively, GS-PAT removes the target points amplitude constraints, combining forward and backward propagation for high computational speed [GSPAT], while Weighted Gerchberg-Saxton (WGS) weights the pressure at each target point during each iteration to achieve low variance in pressure across target points[TGS]. Nevertheless, beyond interpretability, alternating projection algorithms also lack gradient descent’s adaptability to adapt to custom objectives that may be required by users for different applications.
Gor’kov potential, Force, and Stiffness: When the field has been modelled for a given set of transducer activations, various metrics can be computed that quantify the field. The simplest measure of an acoustic field is the pressure at a point in space. When the solver has correctly focused waves at a point, there should be a local maxima of pressure at that point and the amplitude of the sound field at position can be described by $|\boldsymbol{A}\boldsymbol{x}|$. Similarly, the phase can be found by taking the complex argument of $\boldsymbol{A}\boldsymbol{x}$.
The Gor’kov potential is a potential field based on both the pressure and gradient of pressure surrounding a small particle as [AcousticElements]
@@
Here, $c_{0}$ and $c_{p}$ are the speed of sound in the medium and the particle, $\rho_{0}$ and $\rho_{p}$ are the density of the medium and particle, $\omega$ is the angular frequency, and $V$ is the volume of the particle. $p$ is the pressure at a point and ${\partial p}/{\partial a}$ is the gradient of pressure in the $a$-axis where $a\in(x,y,z)$ [AcousticElements]. The Gor’kov potential is the field such that the negative gradient of the Gor’kov potential is equal to the force acting on a small particle in the sound field.
$\boldsymbol{F}_{rad}=-\boldsymbol{\nabla}U.$ (11)
As a result, the point at which a particle will levitate corresponds to where the gradient of Gor’kov potential is zero (so there is no net force at that point). In reality, the levitation location will be very slightly below this point due to the force of gravity but this difference can be largely neglected for very small particles [ReviewOfProgress]. The points with $\boldsymbol{\nabla}U=0$ includes both minima and maxima of Gor’kov potential and both are technically valid levitation locations (with net force equal to zero). However, at a maxima of Gor’kov potential, the forces diverge, and therefore the trap becomes unstable. To quantify a stable trap, we can use the stiffness at a point in space, which can be represented as the sum of the unmixed, second gradients (the Laplacian) of the Gor’kov potential as [AcousticElements]
@@
2.2 Existing Software
AcousTools OpenMPD[OpenMPD] Ultraino[Ultraino] Levitate[Levitate]